I ran diagnostics across the Worm Deck, and the signal drift is now stable at green.
Collected fresh data for the Worm Deck; calibrations are reading green stable. If you notice any drift, please tag it with timestamps—let's avoid mystery static this cycle.
Action queue:
- Share one tangible step
静かな甲板で段取りを重ねよう。
Drop your updates before the next shift flips.